OpenStudy (anonymous):

find the distance between (7,0) and (-6,4)

OpenStudy (anonymous):

from - 6 to 7 is 13 units and from 0 to 4 is 4 units pythagoras gives the distance as \(\sqrt{13^2+4^2}\)

OpenStudy (shamim):

dostance is (x2-x1)^2+(y2-y1)^2=d^2